OpenStudy (nevillesnetwork):

Medal & Fan (Please help asap! Im behind on 32 assignments because of family events) HELP PLZ! The label on the car's antifreeze container claims to protect the car between −40°C and 125°C. To convert Celsius temperature to Fahrenheit temperature, the formula is C = 5/9 (F − 32). Write and solve the inequality to determine the Fahrenheit temperature range at which this antifreeze protects the car. (2 points) −40 > 5/9 (F − 32) > 125; −40 > F > 257 −40 < 5/9 (F − 32) < 125; −40 < F < 257 −40 < 5/9 (F − 32); −40 < F 5/9 (F − 32) < 125; F < 257

OpenStudy (nevillesnetwork):

@triciaal

OpenStudy (aveline):

You need to change your given equation so that you can find the degrees fahrenheit. C = 5/9 (F − 32) F = 9/5C +32 So... ((9/5 x -40°C) +32) < °F < ((9/5 x 125°C) + 32)

OpenStudy (aveline):

You'll need to simplify the right side of the inequality to find your answer.

OpenStudy (nevillesnetwork):

This right side, simplify this? > 9/5 x 125°C) + 32 @Aveline

OpenStudy (aveline):

Yes.

OpenStudy (nevillesnetwork):

\[9/5x ^{125} + 32\]

OpenStudy (nevillesnetwork):

Is this the answer ^ @Aveline

OpenStudy (triciaal):